1. The electric potential (V) in a region of space in volts is given by

V=5x-7x2y+8y2+16yz-5z , where distances are measured in meter. Calculate the y-component of the field at point (2 4 -3).


1
Expert's answer
2021-10-11T07:01:12-0400

 The y-component of ekectric field is given by

"E_y=-\frac{\partial V}{\partial y}=-\frac{\partial }{\partial y}(5x-7x^2y+8y^2+16yz-5z)"

"E_y=7x^2-16y-16z\\
=7*2^2-16*4-16*(-3)=12\:\rm V/m"